OpenStudy (anonymous):

Andy has two rectangular gardens. His smaller garden is 8 feet long and 4 feet wide. The larger garden is 12 feet long and six feet wide. Which statement correctly describes Andy gardens?

OpenStudy (anonymous):

A) The gardens are only congruent. B) The gardens are similar and congruent. Eliminate C) The gardens are congruent, but not similar. D) The gardens are similar, but not congruent.

OpenStudy (anonymous):

Well, if we write a proportion for each of the gardens:\[\frac{ 4 }{ 8 }\] and \[\frac{ 6 }{ 12 }\] They can both be simplified down to 1/2. They are clearly not congruent since they are different lengths.

OpenStudy (anonymous):

Does that mean the answers D?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Thank You! :)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

@AnonymousBeast is correct that they certainly are not congruent. If the two garden were congruent, the individual measurements for length and width would be exactly the same. However, I propose setting up proportions of the length of Garden A to length of Garden B and the same for the width... ie. \[\frac{ 8 }{ 12 } = \] and \[\frac{ 4 }{ 6 }=\] If you reduce them both, you get \[\frac{ 2 }{ 3 } and \frac{ 2 }{ 3 }\]. These are EQUAL PROPORTIONS. Therefore, because they are proportionate to one another, the two gardens are SIMILAR.

OpenStudy (anonymous):

If there were an additional answer choice of Neither congruent, nor similar, you would have to do this step in order to determine whether or not they are actually similar.

OpenStudy (anonymous):

Oh ok thanks for the response :)

OpenStudy (anonymous):

@bcgfghg You're welcome.

OpenStudy (anonymous):

Wanna answer this one too i think its right. Which type of triangle has side lengths of 2 in, 5 in, and 6 in? A) acute Eliminate B) equilateral C) obtuse D) right @TOMWALK @AnonymousBeast

OpenStudy (anonymous):

Well, it cannot be equilateral because the sides are not the same length. Nor can it be a right triangle because: According to the Pythagorean Theorem \[2^2+5^2=6^2\]But, it does not because \[29\neq36\] So B and D are eliminated.

OpenStudy (anonymous):

In order for a triangle to be a right triangle, the Pythagorean theorem needs to apply. In other words: The sum of the squares of the two shorter sides must equal the square of the longest side, which is the hypotenuse.

OpenStudy (anonymous):

@AnonymousBeast is correct.

OpenStudy (anonymous):

The answer was obtuse which is what I picked based on your guys help. Thanks

OpenStudy (anonymous):

Yes it was, my apologies, it was taking me a bit of time to type.